# Nuisance-free Automatic Ground Collision Avoidance System Design: Merging Exponential-CBF and Adaptive Sliding Manifolds

### v1

@@ V-A1 Scenario-1: Low Altitude
Figure 5: Corresponding state trajectories during recovery maneuver: “pilot” subscript denotes the pilot command, “GCAS” subscript denotes the Auto-GCAS command, “cmd” subscript denotes the overall resultant command, and “A/C” subscript denotes the aircraft response.
@@
Figure 6: Corresponding control surface deflections and control allocation objective, $||\phi u-\tau_{c}||$.
@@
V-A2 Scenario-2: High Altitude
@@
Figure 8: Corresponding state trajectories during recovery maneuver: “pilot” subscript denotes the pilot command, “GCAS” subscript denotes the Auto-GCAS command, “cmd” subscript denotes the overall resultant command, and “A/C” subscript denotes the aircraft response.
@@
Figure 9: Corresponding control surface deflections and control allocation objective, $||\phi u-\tau_{c}||$.
@@
V-B Authority-sharing Assessment
@@
Figure 11: Corresponding state trajectories during recovery maneuver: “pilot” subscript denotes the pilot command, “GCAS” subscript denotes the Auto-GCAS command, “cmd” subscript denotes the overall resultant command, and “A/C” subscript denotes the aircraft response.
@@
Figure 12: Corresponding control surface deflections and control allocation objective, $||\phi u-\tau_{c}||$.
@@
V-C Monte Carlo Simulations
Monte Carlo simulations are conducted under 500 randomized initial conditions to investigate the efficacy of the proposed method. In this assessment, since the DTED is assumed to be flat and at ground level (0 meters), the previously mentioned terrain scan pattern has not been utilized. All the states of the aircraft are randomized, as depicted in Fig. 14.
@@ V-C Monte Carlo Simulations
[…]
@@
[…]
VI Conclusion
@@
The efficacy of the overall proposed framework has been evaluated through three distinct scenarios and Monte Carlo simulations, conducted under $500$ randomized different initial conditions. In terrain collision avoidance assessments, timely and aggressive commands are generated by the Auto-GCAS, providing maximum amplitude pitch rate commands and maintaining sufficiently close proximity to the buffer zone. Additionally, an authority-sharing scenario is examined, with different pilot commands tested at various points in time. Pilot input is allowed by the Auto-GCAS as long as the imposed constraints are not violated, ensuring that unnecessary interventions do not occur. Finally, Monte Carlo simulations, including $500$ different cases, demonstrate that the proposed framework is highly effective in protecting the aircraft from ground collisions while generating timely and aggressive commands. The success is indicated by the maximum amplitude of commands, load factor, and Euclidean distance between the generated commands and the reference nuisance-free command. The proposed system achieved a remarkable success rate of $99.8\%$ across $500$ randomized Monte Carlo simulations. These results highlight the potential of the proposed framework to significantly mitigate controlled flight into terrain incidents, fostering trust and collaboration between pilots and automation systems.
Future work will focus on refining the optimization of the ECBF control gain vector in relation to dynamic terrain scan patterns. Additionally, implementing and testing the framework in real-time on a fixed-wing UAV test-bed will further validate its operational feasibility and scalability.

### v2

@@ 5.1.1 Scenario-1: Low Altitude
Figure 5: Corresponding state trajectories: “pilot” = pilot input, “GCAS” = Auto-GCAS command, “cmd” = resultant command, “A/C” = aircraft response.
@@
Figure 6: Corresponding control surface deflections, control allocation objective, $\phi u-\tau_{c}||$, and adaptive ECBF control gain vector.
@@
5.1.2 Scenario-2: High Altitude
@@
Figure 8: Corresponding state trajectories: “pilot” = pilot input, “GCAS” = Auto-GCAS command, “cmd” = resultant command, “A/C” = aircraft response.
@@
Figure 9: Corresponding control surface deflections, control allocation objective, $\phi u-\tau_{c}||$, and adaptive ECBF control gain vector
@@
5.2 Authority-sharing Assessment
@@
Figure 11: Corresponding state trajectories: “pilot” = pilot input, “GCAS” = Auto-GCAS command, “cmd” = resultant command, “A/C” = aircraft response.
@@
Figure 12: Corresponding control surface deflections, control allocation objective, $\phi u-\tau_{c}||$, and adaptive ECBF control gain vector
@@
5.3 Monte Carlo Simulations
Monte Carlo simulations are conducted under $850$ randomized initial conditions to investigate the efficacy of the proposed method. While randomized initial conditions are used to evaluate general robustness, additional tailored scenarios—such as steep dives and low-altitude banked turns—has already been presented to challenge the controller under known high-risk conditions. In this assessment, since the DTED is assumed to be flat and at ground level (0 meters), the previously mentioned terrain scan pattern has not been utilized. All the states of the aircraft are randomized, as depicted in Fig. 14.
@@ 5.3 Monte Carlo Simulations
The angle of attack is randomized within the interval $10^{\circ},\hskip 4.26773pt20^{\circ}]$, sideslip within $10^{\circ},\hskip 4.26773pt10^{\circ}]$, bank angle within $150^{\circ},\hskip 4.26773pt150^{\circ}]$, pitch angle within $75^{\circ},\hskip 4.26773pt-10^{\circ}]$, yaw angle within $30^{\circ},\hskip 4.26773pt30^{\circ}]$, roll rate within $120^{\circ}/\text{s},\hskip 4.26773pt120^{\circ}/\text{s}]$, pitch rate within $10^{\circ}/\text{s},\hskip 4.26773pt20^{\circ}/\text{s}]$, yaw rate within $15^{\circ}/\text{s},\hskip 4.26773pt15^{\circ}/\text{s}]$, true airspeed within $200\text{m/s},\hskip 4.26773pt350\text{m/s}]$, and altitude within $1000\text{m},\hskip 4.26773pt5000\text{m}]$. Understandably, the search space is meaningfully broad. The performance metrics for the evaluation are the minimum altitude, maximum pitch rate command, maximum load factor, and the Euclidean distance between the reference command depicted in Fig. 1 and the applied pitch rate command. In order to obtain a meaningful Euclidean distance metric, the applied pitch rate and reference commands are normalized between 0 and $1$. Then, the dynamic time warping method is employed to compare the similarity between the shapes of the signals by aligning them in time to minimize differences, even if they are out of phase or vary in speed. This is achieved using MATLAB®’s dtw function, which calculates an optimal match between two sequences by warping their time axes to minimize the cumulative distance between corresponding points. Consequently, the comparison is independent of the application time and magnitudes. The statistical illustration of the Monte Carlo simulations are presented in Fig. 15.
@@
The results indicate a high success rate of the proposed methodology. The average minimum altitude is 100 meters, which corresponds to the predetermined buffer height. This outcome is a clear indication of timely intervention. Additionally, the maximum pitch rate command and maximum load factor reflect the aggressiveness of the Auto-GCAS intervention. The outcomes that should be elaborated are the failed case and the Euclidean distance. The initial conditions for the failed case are $1120.6$ meters for the altitude, $64.86^{\circ}$ for the pitch angle, $96.96^{\circ}$ for the bank angle, and $330.1\,\text{m/s}$ for the true airspeed. These parameters are of concern due to their significant impact in a dive scenario. Notably, the altitude is relatively low to protect the aircraft from such a steep dive ($64.86^{\circ}$ for the pitch angle) and high speed ($330.1\,\text{m/s}$, circa 1 Mach). The histories of this case reveal that the applied pitch rate corresponds to a 9G pull-up at the beginning of the dive and it sustained for approximately 6 seconds; however, this recovery maneuver still fails to prevent a collision. Consequently, this case can be regarded as both extreme and harsh. Furthermore, the Euclidean distance should be as close to 0 as possible for a perfect overlap between the reference signal and the applied pitch rate command. The average Euclidean distance is approximately $5$. To interpret this value, a comparison is essential. Therefore, the Euclidean distances of the pitch rate commands depicted in Fig. 5 and Fig. 8 are $1.2$ and $5.1611$, respectively, and can be regarded as reference values. Thus, it can be interpreted that the applied pitch rate commands closely resemble the pitch rate shown in Fig. 8, which is satisfactory in terms of the nuisance-free criterion. Consequently, it is evident that the proposed methodology achieves its objectives with a high success rate of $849$ successful collision avoidance out of $850$ random cases, indicating a $99.88\%$ success rate.
6 Discussion
This section discusses the practical applicability, comparative advantages, aircraft suitability, and limitations of the proposed nuisance-free Auto-GCAS framework, with emphasis on robustness and implementation feasibility.
6.1 Practical Implementation and Applicability
The proposed method is designed with real-time implementation in mind. By formulating collision avoidance and flight envelope protection as convex quadratic programs, the algorithm can be integrated into modern flight control computers without requiring extensive computational resources. The adaptive ECBF gains are computed offline across a range of flight conditions (pitch angle, bank angle, airspeed), enabling efficient onboard scheduling without online optimization. Integration into existing avionics architectures is feasible, requiring access to standard state estimates, DTED, and pilot commands, which are common in current modern and advanced aircraft regardless of manned or unmanned. Additionally, the framework is applicable to fixed-wing aircraft platforms, including manned and unmanned aircraft with comparable control authority. It is particularly suitable for avoiding terrain-based obstacles represented with DTED, such as hills or mountainous terrain.
6.2 Addressing Literature Gaps and Comparative Strengths
Unlike prior nuisance-free Auto-GCAS approaches that rely on computationally demanding trajectory prediction or simplified 3DoF dynamics, this method uses adaptive ECBFs to generate timely and aggressive recovery maneuvers without sacrificing model fidelity. Additionally, it incorporates a flight envelope protection layer based on control barrier functions, addressing a key shortcoming in earlier studies that lacked comprehensive safety enforcement. Simulation results, including Monte Carlo assessments with a $99.88\%$ success rate, demonstrate the method’s ability to meet all critical objectives: ground collision avoidance, nuisance mitigation, and envelope protection. The avoidance behavior is both continuous and aggressive, closely emulating pilot-preferred interventions and minimizing false alarms.
6.3 Limitations and Robustness Considerations
Despite its strengths, the method has some limitations. While the gain schedule improves adaptivity, its resolution may limit performance in rapidly changing flight conditions. This could be addressed by incorporating real-time interpolation or machine learning-based gain tuning. The symmetric flight assumption (due to the bank-to-level requirement), although practical, restricts applicability in highly asymmetric or failure scenarios. Additionally, the fixed terrain scan pattern could be enhanced with adaptive or predictive scanning for improved situational awareness. Furthermore, performance over real geographic terrain requires further investigation, as the current assessment has been conducted using a simple synthetic terrain.
In terms of robustness, the CBF-based formulation offers inherent resilience to bounded model uncertainties. However, the method still relies on nominal aerodynamic models, and its performance could degrade under severe discrepancies. Future work may explore robust or uncertainty-aware CBF formulations to maintain safety guarantees across wider operational envelopes.
7 Conclusion
@@
The efficacy of the overall proposed framework has been evaluated through three distinct scenarios and Monte Carlo simulations, conducted under $850$ randomized different initial conditions. In terrain collision avoidance assessments, timely and aggressive commands are generated by the Auto-GCAS, providing maximum amplitude pitch rate commands and maintaining sufficiently close proximity to the buffer zone. Additionally, an authority-sharing scenario is examined, with different pilot commands tested at various points in time. Pilot input is allowed by the Auto-GCAS as long as the imposed constraints are not violated, ensuring that unnecessary interventions do not occur. Finally, Monte Carlo simulations, including $850$ different cases, demonstrate that the proposed framework is highly effective in protecting the aircraft from ground collisions while generating timely and aggressive commands. The success is indicated by the maximum amplitude of commands, load factor, and Euclidean distance between the generated commands and the reference nuisance-free command. The proposed system achieved a remarkable success rate of $99.88\%$ across $850$ randomized Monte Carlo simulations. These results highlight the potential of the proposed framework to significantly mitigate controlled flight into terrain incidents, fostering trust and collaboration between pilots and automation systems.
Consequently, this work contributes to the broader body of safety-critical flight control by advancing the application of adaptive and constraint-aware CBF-based methods to the Auto-GCAS domain. The results offer actionable insights for researchers aiming to develop pilot-compatible, certifiable avoidance systems. Future work will focus on refining the optimization of the ECBF control gain vector in relation to dynamic terrain scan patterns. Additionally, implementing and testing the framework in real-time on a fixed-wing UAV test-bed will further validate its operational feasibility and scalability.
Acknowledgments
During the preparation of this work the author(s) used ChatGPT in order to improve language and readability. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication. Also, the authors are deeply grateful to Dr. Ersin Daş for his valuable discussions during the presentation of this study.